A farmer bought a number of pigs for $252. However, 7 of them died before he could sell the rest at a profit of 9 per pig. His total profit was $126. How many pigs did he originally buy?

Found 2 solutions by stanbon, JoelSchwartz:
Answer by stanbon(75887)

A farmer bought a number of pigs for $252. However, 7 of them died before he could sell the rest at a profit of $9 per pig. His total profit was $126. How many pigs did he originally buy?
--------
# he originally bought:: "x"
Price per pig 252/x
-------------------------------------
After 7 died
# of pigs:: x-7
Profit on the x-7:: 9x-63
----
Equation:
9x-63 = 126
9x = 189
x = 21 (original # of pigs)
--------------
Cheers,
Stan H.
--------------

Answer by JoelSchwartz(130)

x=cost of each pig
y=the number of pigs bought
xy=252
x=252/y
y-7=the new number of pigs when the farmer sells the pigs
x+9=the price of each pig when the farmer sells the pig for nine dollars more than he paid for each pig
(y-7)(x+9)=252+126
(y-7)(x+9)=378
yx-7x+9y-63=378
252-7x+9y-63=378
9y-7x+189=378
9y-7x=189
9y-7(252/y)=189
(9y^2)-7*252=189y
(9y^2)-189y-1764=0
(y^2)-196-21y=0
((y-10.5)^2)=(y^2)-21y+110.25
((y-10.5)^2)-306.25=0
((y-10.5)^2)=306.25
y-10.5=+-17.5
y=-7 and y=28
x=252/28
x=9
y=28
He originally bought twenty eight pigs